# TECP: Token-Entropy Conformal Prediction for LLMs

### v1

@@ Abstract
Uncertainty quantification (UQ) for open-ended language generation remains a critical yet underexplored challenge, especially under black-box constraints where internal model signals are inaccessible. In this paper, we introduce Token-Entropy Conformal Prediction (TECP), a novel framework that leverages token-level entropy as a logit-free, reference-free uncertainty measure and integrates it into a split conformal prediction (CP) pipeline to construct prediction sets with formal coverage guarantees. Unlike existing approaches that rely on semantic consistency heuristics or white-box features, TECP directly estimates epistemic uncertainty from the token entropy structure of sampled generations and calibrates uncertainty thresholds via CP quantiles to ensure provable error control. Empirical evaluations across six large language models and two benchmarks (CoQA and TriviaQA) demonstrate that TECP consistently achieves reliable coverage and compact prediction sets, outperforming prior self-consistency–based UQ methods. Our method provides a principled and efficient solution for trustworthy generation in black-box LLM settings.
@@ 1 Introduction
Despite their strong performance across many tasks, LLMs still exhibit pervasive reliability issues (Jiang et al., ; Huang et al., 2025; Lavrinovics et al., 2025; Zhou et al., 2025), including hallucinations and factual errors, and may produce responses that appear well-formed yet are fabricated or detached from reality, thereby weakening their suitability for deployment in high-stakes settings. Against this backdrop, uncertainty quantification (UQ) has become a key approach for assessing the credibility of model outputs, enabling users to determine when a model’s answers can be trusted (Mora-Cross & Calderon-Ramirez, 2024; Farquhar et al., 2024; Duan et al., 2024; Qiu & Miikkulainen, 2024; Wang et al., 2024, 2025f). This is especially critical in high-risk professional domains such as medicine and psychology, where decision-making often relies heavily on LLM-generated content. However, uncertainty arises from heterogeneous sources, including epistemic and aleatoric components, so developing effective and well-grounded approaches to quantifying uncertainty remains an urgent and important issue (Wang et al., 2025a, d, c; Chen et al., 2024c).
@@
Compared to traditional probabilistic prediction, CP imposes minimal assumptions on data distribution—requiring only that samples satisfy the property of exchangeability—which endows it with strong adaptability and robustness across diverse tasks and model architectures. Moreover, the size of the prediction set directly reflects model uncertainty: when confidence is low, the set expands to preserve coverage; when confidence is high, the set contracts to enhance decision precision. Owing to its provable coverage guarantee, CP is of particular importance in high-stakes applications such as medical diagnosis, legal text analysis, and scientific research, as it not only quantifies the reliability of predictions but also enables flexible trade-offs between risk and utility under varying confidence thresholds. In recent years, researchers have further extended CP beyond traditional classification and regression tasks to more complex domains such as open-ended natural language generation and structured prediction. By incorporating techniques such as sampling and reprompting, these advancements effectively address the challenges posed by vast output spaces and the inherent diversity of possible answers.
We undertake a systematic empirical investigation of six cutting-edge large language models—LLaMA-3.2-1B, Qwen2.5-3B-Instruct, Vicuna-7B-v1.5, Qwen2.5-7B-Instruct, LLaMA-3.1-8B-Instruct, and Vicuna-13B-v1.5—across the CoQA and TriviaQA benchmarks, with the objective of rigorously evaluating their uncertainty quantification capabilities. The assessment is grounded in two principled metrics: (1) Empirical Coverage Rate (ECR), quantifying the proportion of instances wherein the constructed prediction set encapsulates the ground-truth response, and (2) Average Prediction Set Size (APSS), measuring the expected cardinality of prediction sets as a proxy for epistemic uncertainty. Experimental analyses uncover pronounced disparities in ECR and APSS across models and datasets, underscoring the sensitivity of uncertainty estimates to both model-specific generative competence and dataset-inherent statistical properties. To address the inherent challenges of uncertainty estimation in open-ended generation, we introduce a token-level entropy-based nonconformity scoring paradigm within the conformal prediction framework. Our method enables the construction of calibrated prediction sets under minimal distributional assumptions and without reliance on labeled supervision. Empirical results substantiate the efficacy of the proposed approach in maintaining rigorous error control while achieving high coverage of ground-truth outputs, thereby furnishing robust reliability guarantees for model predictions in open-domain generative settings.
@@ 2 Related Work
Uncertainty Analysis: Framed by uncertainty analysis, prior work traces a progression from general UQ to black-box UQ for LLMs and, ultimately, to CP-based coverage guarantees: in ML and NLP, UQ is foundational for decision-making and risk control, yet confidence-based metrics (e.g., entropy) are vulnerable to calibration mismatch, while Bayesian and ensemble paradigms—despite their theoretical appeal—incur prohibitive computational and engineering overheads at LLM scale. Early LLM efforts relied on white-box signals (token likelihoods, internal activations) and fine-tuning for improved calibration, but API-level black-box constraints and resource budgets limit their practicality, motivating an output-only, black-box route: multi-sample generation is used to capture semantic consistency and dispersion, with the “most frequent generation” serving as a self-consistency anchor to construct actionable, internal-agnostic uncertainty surrogates. Building on this, conformal prediction (CP) offers a systematic bridge from heuristic scores to coverage-guaranteed prediction sets by using a small number of i.i.d. calibration instances to map external UQ scores to set thresholds and to achieve ground-truth coverage at a user-specified error rate, while remaining model- and distribution-agnostic and computationally feasible. Further, to accommodate the vast output space of open-ended generation, recent work extends CP from classification to sequence generation via stopping rules and sampling approximations, thereby making explicit both “when confidence is sufficient” and “how large the set must be to remain robust” (Iutzeler & Mazoyer, 2025).
@@ 3 Method
This section proposes a general framework grounded in semantic self-consistency and split conformal prediction to assess uncertainty in outputs generated by black-box language models and, on this basis, to construct prediction sets endowed with statistical coverage guarantees. Unlike traditional approaches that rely on log-likelihoods or internal distributions, we treat the language model as a complete black box: only standard API access to generated outputs is assumed, with no visibility into parameters, logits, or attention weights. Under this setting, the central challenge is to estimate confidence from a finite set of sampled outputs and to achieve theoretically verifiable calibrated prediction.
@@ 3.1 Problem Setup and Candidate Generation Mechanism
$\hat{y}_{m}\sim f(x;\theta),\quad m=1,\ldots,M,$ (1)
@@
Figure 1: EMR vs. Alpha for six models on TriviaQA dataset using TECP
@@ 3.2 Uncertainty Estimation Driven by Inter-Candidate Semantic Consistency
where $L_{m}$ is the length of $\hat{y}_{m}$, $\mathcal{V}$ is the vocabulary, and $p_{t}(v)$ is the model’s predictive distribution over tokens $v$ at position $t$. Higher cumulative entropy indicates greater uncertainty and, hence, lower confidence in the generation. We adopt token-level entropy as our uncertainty metric because it captures the model’s intrinsic uncertainty during generation, avoiding the overconfidence bias and hallucination sensitivity inherent in frequency-based diversity measures.This reference-free criterion evaluates each candidate’s confidence independently of ground truth.
@@ 3.3 Uncertainty Calibration and Prediction-Set Construction
    Data partitioning and filtering. For each sample $(x_{i},y_{i}^{*})$, if there exists a candidate $\hat{y}_{m}$ such that $S(\hat{y}_{m},y_{i}^{*})\geq\tau$, the sample is deemed assessable (i.e., of acceptable generation quality) and retained in $\mathcal{D}_{\text{filtered}}$, where $\tau$ is a semantic-matching threshold typically in the range 0.80.80.90.9. We then randomly split $\mathcal{D}_{\text{filtered}}$ into a calibration subset $\mathcal{D}_{\text{cal}}$ and a test subset $\mathcal{D}_{\text{test}}$ in a user-specified ratio.
@@
Figure 2: EMR vs. Alpha for six models on TriviaQA dataset using ConU
@@ 4.1 Experimental Setup
Backbone LLMs and Evaluation Tasks. Since we incorporate token entropy into the conformal prediction framework for open-ended Question Answering (QA) tasks, it is necessary to rigorously evaluate the performance of the proposed architecture. To this end, we select a diverse set of open-source large language models (LLMs) as backbone models, including the Llama series (Llama-3.2-1B, Llama-3.1-8B-Instruct), the Qwen series (Qwen2.5-3B-Instruct, Qwen2.5-7B-Instruct), and the Vicuna series (Vicuna-7B-v1.5, Vicuna-13B-v1.5).
Datasets. We employ TriviaQA and CoQA as our experimental datasets. The TriviaQA dataset contains over 650,000 question–answer–evidence triples, covering a wide range of topics such as history, science, and entertainment, and supports multiple tasks including open-domain question answering and extractive question answering. CoQA, is a large-scale dataset designed for building conversational question answering systems.
@@
Figure 3: EMR vs. Alpha for six models on CoQA dataset using TECP
@@
  • ConU. We propose a prediction method based on Conformal Uncertainty, which evaluates candidate responses generated for each input prompt and quantifies predictive uncertainty by incorporating both model confidence scores and calibration error. In contrast to heuristic approaches that rely solely on response-level variability, ConU offers a principled mechanism for constructing prediction sets with theoretical risk control guarantees. Notably, the method does not require access to model log-probabilities, yet is still able to selectively retain high-confidence responses. This enables the construction of prediction sets that are statistically guaranteed to meet predefined risk levels, while remaining representative, stable, and interpretable.
@@ 4.2 Results for QA
Figure 4: EMR vs. Alpha for six models on CoQA dataset using ConU
@@ 4.3 Ablation Study
To systematically analyze the effect of different split ratios on prediction set coverage performance, we conduct an ablation study on two open-domain question answering datasets: TriviaQA and CoQA. In this experiment, we vary the calibration–test split ratios across three configurations: 0.3, 0.5, and 0.7, corresponding to different proportions of calibration data. All evaluations are performed under a fixed risk level of $\alpha=0.1$, and the results are averaged over 100 random seeds to mitigate the influence of sampling variability. The resulting empirical coverage (defined as $1-\text{EMR}$) is visualized using line charts in Figures 1 and 3, which illustrate model-wise performance under each split ratio.
@@ 5 Conclusion
Our approach reflects a typical limitation of white-box conformal prediction frameworks—namely, the reliance on uncertainty scores derived from model output probabilities. While predictive uncertainty facilitates the construction of prediction sets with formal coverage guarantees, the effectiveness of calibration is intrinsically linked to the accuracy and stability of the underlying uncertainty signal, which may be compromised by instruction tuning or reinforcement learning. Furthermore, the method presumes that such uncertainty scores can be reliably extracted across diverse models and tasks, an assumption that may break down under distribution shifts or non-standard decoding protocols. Future directions include developing scoring functions that are more robust to model-specific artifacts, or exploring hybrid frameworks that integrate both internal and external uncertainty cues to improve generalization.

### v2

@@ Abstract
Uncertainty quantification (UQ) for open-ended language generation remains a critical yet underexplored challenge, especially under black-box constraints where internal model signals are accessible. In this paper, we introduce Token- Entropy Conformal Prediction (TECP). This novel framework leverages token-level entropy as a logit-free, reference-free uncertainty measure and integrates it into a split conformal prediction (CP) pipeline to construct prediction sets with formal coverage guarantees. Unlike existing approaches that rely on semantic consistency heuristics or white-box features, TECP directly estimates episodic uncertainty from the token entropy structure of sampled generations and calibrates uncertainty thresholds via CP quantiles to ensure provable error control. Empirical evaluations across six large language models and two benchmarks (CoQA and TriviaQA) demonstrate that TECP consistently achieves reliable coverage and compact prediction sets, outperforming prior self UQ methods. Our method provides a principled and efficient solution for trustworthy generation in black-box LLM settings.
@@ 1 Introduction
Despite their strong performance across many tasks, LLMs still exhibit pervasive reliability issues (Jiang et al., ; Huang et al., 2025; Lavrinovics et al., 2025; Zhou et al., 2025), including hallucinations and factual errors, and may produce responses that appear well-formed yet are fabricated or detached from reality. It reduces their adaptability for deployment in high-stakes scenarios. Against this backdrop, uncertainty quantification (UQ) has become a key approach for assessing the credibility of model outputs, enabling users to determine when a model’s answers can be trusted (Mora-Cross & Calderon-Ramirez, 2024; Farquhar et al., 2024; Duan et al., 2024; Qiu & Miikkulainen, 2024; Wang et al., 2024, 2025f). It is especially critical in high-risk professional domains such as medicine and psychology, where decision-making often relies heavily on LLM-generated content. However, uncertainty arises from heterogeneous sources, including epistemic and aleatoric components, so developing effective and well-grounded approaches to quantifying uncertainty remains an urgent and important issue (Wang et al., 2025a, d, c; Chen et al., 2024c).
@@
Compared to traditional probabilistic prediction, CP imposes minimal assumptions on data distribution, requiring only that samples satisfy the property of exchangeability—which endows it with strong adaptability and robustness across diverse tasks and model architectures. Moreover, the size of the prediction set can directly indicate the model’s uncertainty: when confidence is low, the set expands to preserve coverage; when confidence is high, the set contracts to enhance decision precision. Owing to its provable coverage guarantee, CP is of particular importance in high-stakes applications such as medical diagnosis, legal text analysis, and scientific research, as it not only quantifies the reliability of predictions but also enables flexible trade-offs between risk and utility under varying confidence thresholds. In recent years, researchers have further extended CP beyond traditional classification and regression tasks to more complex domains such as open-ended natural language generation and structured prediction. By incorporating techniques such as sampling and reprompting, these advancements effectively address the challenges posed by vast output spaces and the inherent diversity of possible answers.
We undertake a systematic empirical investigation of six cutting-edge large language models—LLaMA-3.2-1B, Qwen2.5-3B-Instruct, Vicuna-7B-v1.5, Qwen2.5-7B-Instruct, LLaMA-3.1-8B-Instruct, and Vicuna-13B-v1.5—across the CoQA and TriviaQA benchmarks, with the objective of rigorously evaluating their uncertainty quantification capabilities. The assessment is grounded in two principled metrics: (1) Empirical Coverage Rate (ECR), quantifying the proportion of instances wherein the constructed prediction set encapsulates the ground-truth response, and (2) Average Prediction Set Size (APSS), measuring the expected cardinality of prediction sets as a proxy for epistemic uncertainty. Experimental analyses uncover pronounced disparities in ECR and APSS across models and datasets, underscoring the sensitivity of uncertainty estimates to both model-specific generative competence and dataset-inherent statistical properties. To address the inherent challenges of uncertainty estimation in open-ended generation, we introduce a token-level entropy-based nonconformity scoring paradigm within the conformal prediction framework. The proposed method enables the construction of calibrated prediction sets under minimal distributional assumptions and without reliance on labeled supervision. Empirical results substantiate the efficacy of the proposed approach in maintaining rigorous error control while achieving high coverage of ground-truth outputs, thereby furnishing robust reliability guarantees for model predictions in open-domain generative settings.
@@ 2 Related Work
Uncertainty Analysis: Framed by uncertainty analysis, prior work traces a progression from general UQ to black-box UQ for LLMs and, ultimately, to CP-based coverage guarantees: in ML and NLP, UQ is foundational for decision-making and risk control, yet confidence-based metrics (e.g., entropy) are vulnerable to calibration mismatch, while Bayesian and ensemble paradigms—despite their theoretical appeal—incur prohibitive computational and engineering overheads at LLM scale. Early LLM efforts relied on white-box signals (token likelihoods, internal activations) and fine-tuning for improved calibration. However, API-level black-box constraints and resource budgets limit their practicality, motivating an output-only, black-box route: multi-sample generation is used to capture semantic consistency and dispersion, with the “most frequent generation” serving as a self-consistency anchor to construct actionable, internal-agnostic uncertainty surrogates. Building on this, conformal prediction (CP) offers a systematic bridge from heuristic scores to coverage-guaranteed prediction sets by using a small number of i.i.d. calibration instances to map external UQ scores to set thresholds and to achieve ground-truth coverage at a user-specified error rate, while remaining model- and distribution-agnostic and computationally feasible. Further, to accommodate the vast output space of open-ended generation, recent work extends CP from classification to sequence generation via stopping rules and sampling approximations, thereby making explicit both “when confidence is sufficient” and “how large the set should be to remain robust” (Iutzeler & Mazoyer, 2025).
@@ 3 Method
This section proposes a general framework grounded in semantic self-consistency and split conformal prediction to assess uncertainty in outputs generated by black-box language models and, on this basis, to construct prediction sets endowed with statistical coverage guarantees. Unlike traditional approaches that rely on log-likelihoods or internal distributions, we treat the language model as a complete black box: only standard API access to generated outputs is assumed, with no visibility into parameters, logits, or attention weights. Under this setting, the central challenge is to estimate confidence from a finite set of sampled outputs and to achieve a theoretically verifiable calibrated prediction.
@@ 3.1 Problem Setup and Candidate Generation Mechanism
$\hat{y}_{m}\sim f(x;\theta),\quad m=1,\ldots,M.$ (1)
@@
Figure 1: EMR vs. Alpha for six models on the TriviaQA dataset using TECP
@@ 3.2 Uncertainty Estimation Driven by Inter-Candidate Semantic Consistency
Where $L_{m}$ is the length of $\hat{y}_{m}$, $\mathcal{V}$ is the vocabulary, and $p_{t}(v)$ is the model’s predictive distribution over tokens $v$ at position $t$. Higher cumulative entropy indicates greater uncertainty and, hence, lower confidence in the generation. We adopt token-level entropy as our uncertainty metric because it captures the model’s intrinsic uncertainty during generation, avoiding the overconfidence bias and hallucination sensitivity inherent in frequency-based diversity measures.This reference-free criterion evaluates each candidate’s confidence independently of the ground truth.
@@ 3.3 Uncertainty Calibration and Prediction-Set Construction
    Data partitioning and filtering. For each sample $(x_{i},y_{i}^{*})$, if there exists a candidate $\hat{y}_{m}$ such that $S(\hat{y}_{m},y_{i}^{*})\geq\tau$, the sample is deemed assessable (i.e., of acceptable generation quality) Moreover, retained in $\mathcal{D}_{\text{filtered}}$, where $\tau$ is a semantic-matching threshold typically in the range 0.80.80.90.9. We then randomly split $\mathcal{D}_{\text{filtered}}$ into a calibration subset $\mathcal{D}_{\text{cal}}$ and a test subset $\mathcal{D}_{\text{test}}$ in a user-specified ratio.
@@
Figure 2: EMR vs. Alpha for six models on the TriviaQA dataset using ConU
@@ 4.1 Experimental Setup
Backbone LLMs and Evaluation Tasks. Since we incorporate token entropy into the conformal prediction framework for open-ended Question Answering (QA) tasks, it is necessary to evaluate the performance of the proposed architecture rigorously. To this end, we select a diverse set of open-source large language models (LLMs) as backbone models, including the Llama series (Llama-3.2-1B, Llama-3.1-8B-Instruct), the Qwen series (Qwen2.5-3B-Instruct, Qwen2.5-7B-Instruct), and the Vicuna series (Vicuna-7B-v1.5, Vicuna-13B-v1.5).
Datasets. We employ TriviaQA and CoQA as our experimental datasets. The TriviaQA dataset contains over 650,000 question–answer–evidence triples, covering a wide range of topics such as history, science, and entertainment, and supports multiple tasks including open-domain Question Answering and extractive Question Answering. CoQA, The dataset is large-scale and intended to facilitate the development and construction of conversational Question Answering systems.
@@
Figure 3: EMR vs. Alpha for six models on the CoQA dataset using TECP
@@
  • ConU. A prediction method is proposed based on Conformal Uncertainty, which evaluates candidate responses generated for each input prompt and quantifies predictive uncertainty by incorporating both model confidence scores and calibration error. In contrast to heuristic approaches that rely solely on response-level variability, ConU offers a principled mechanism for constructing prediction sets with theoretical risk control guarantees. Notably, the method does not require access to model log-probabilities, yet is still able to retain high-confidence responses selectively. It enables the construction of prediction sets that are statistically guaranteed to meet predefined risk levels, while remaining representative, stable, and interpretable.
@@ 4.2 Results for QA
Figure 4: EMR vs. Alpha for six models on the CoQA dataset using ConU
@@ 4.3 Ablation Study
To comprehensively evaluate the impact of different split ratios on the coverage performance of the prediction set, we conduct ablation experiments on two open-domain question answering datasets: TriviaQA and CoQA. In this experiment, we vary the calibration–test split ratios across three configurations: 0.3, 0.5, and 0.7, corresponding to different proportions of calibration data. All evaluations are performed under a fixed risk level of $\alpha=0.1$, and the results are averaged over 100 random seeds to mitigate the influence of sampling variability. The resulting empirical coverage (defined as $1-\text{EMR}$) is visualized using line charts in Figures 1 and 3, which illustrate model-wise performance under each split ratio.
@@ 5 Conclusion
Our approach reflects a typical limitation of white-box conformal prediction frameworks—namely, the reliance on uncertainty scores derived from model output probabilities. Although predictive uncertainty supports the creation of prediction sets with theoretical coverage guarantees, the effectiveness of calibration largely depends on the stability and accuracy of the underlying uncertainty signal, which can be affected by instruction tuning or reinforcement learning. Furthermore, the method presumes that such uncertainty scores can be reliably extracted across diverse models and tasks, an assumption that may break down under distribution shifts or non-standard decoding protocols. Future directions include developing scoring functions that are more robust to model-specific artifacts, or exploring hybrid frameworks that integrate both internal and external uncertainty cues to improve generalization.